# Throwing Vines at the Wall: Structure Learning via Random Search

### v1

@@ 3.3 Vine Structure Confidence Sets
On any given data set, we may want to assess whether the model found by Algorithm 1 is indeed better than the benchmark heuristic of (Dissmann et al.,, 2013). For when this is not the case, we may want to continue with the benchmark for qualitative reasons such as ease of interpretability or easier comparison to previous analyses. And even when the benchmark is outperformed, there may be multiple candidates from the random search that are statistically indistinguishable. We tackle this problem via model confidence sets (MCS) (Hansen et al.,, 2011).
Roughly speaking, an MCS is a subset of the candidate that contains the “best ones” with high probability. Further define the set of optimal models according to expected out-of-sample loss as
$\Theta^{*}=\operatorname*{arg\,min}_{\mathcal{V}\in\{\mathcal{V}_{1},\dots,\mathcal{V}_{M}\}}\mathds{E}[L(\widehat{f}_{\mathcal{V},\mathcal{D}_{\text{train}}},\bm{X})\mid\mathcal{D}_{\text{train}}],$
explicitly allowing for the possibility of multiple ones that are equally good. Note that the target set $\Theta^{*}$ is random. It depends on a random draw of candidate models $\mathcal{V}_{1},\dots,\mathcal{V}_{M}$ as well as the quality of the estimates $\widehat{f}_{\mathcal{V},\mathcal{D}_{\text{train}}}$. The latter dependence reflects that some structures may entail densities that are inherently more difficult to estimate than others, which would hurt performance in downstream tasks.
The MCS, which is defined in Definition 3.1 below for the vine context, is a form of discrete argmin inference requiring that all optimal models are included in the set with high probability.
Definition 3.1 (MCS for vine structures).
[…]
@@
$\mathbb{P}(\Theta^{*}\subseteq\widehat{\Theta}\mid\mathcal{D}_{\text{train}},\mathcal{V}_{1},\dots,\mathcal{V}_{M})\geq 1-\alpha.$
An alternative concept is a marginal MCS requiring any fixed $\mathcal{V}^{*}\in\Theta^{*}$ to be included in the set with high probability. The simultaneous variant given in the definition better suits our needs. Indeed, failing to include equivalent models in the set may lead to wrong conclusion regarding the quality of a benchmark model, and potentially lower performance in model averaging (see Section 3.5).
3.4 A Vine MCS Algorithm with Guarantees
Algorithm 2 MCS for Vine Structures.
@@ 3.4 A Vine MCS Algorithm with Guarantees
3:Follow Steps 1-3 from Algorithm 1.
4:Let $\mathcal{D}_{\text{val}}=(\bm{Z}_{i}^{\text{val}})_{i=1}^{n_{\text{val}}}$ and compute per-instance losses for $1\leq i\leq n_{\text{val}}$ and $1\leq m\leq M$:
$L_{i,m}=L(\widehat{f}_{\mathcal{V}_{m},\mathcal{D}_{\text{train}}},\bm{Z}_{i}^{\text{val}}).$
@@
Several methods for estimating MCSs have been proposed in the literature (Hansen et al.,, 2011; Mogstad et al.,, 2024; Zhang et al.,, 2024; Kim and Ramdas,, 2025). In this work, we use the plug-in method of Kim and Ramdas, (2025) (DA-MCS-plug2 in their paper) because of its computational simplicity and good empirical performance. A full description of their method can be found in Appendix A for completeness, along with a runtime comparison with alternative methods.
Algorithm 2 describes how their MCS method fits into our vine structure learning pipeline. The algorithm has runtime complexity $O(Mnd^{2}+Mn)$, where the first term is from the first three steps (see Algorithm 1 and the discussion in Section 3.2), the second from the MCS procedure DA-MCS-plug2, and we refer to Appendix A for this part of the complexity analysis. It also comes with a strong theoretical guarantee, whose proof is given in Appendix B.
@@
Let $M\in\mathds{N}$ be arbitrary and suppose
$\limsup_{n\to\infty}\max_{1\leq m\leq M}\mathds{E}[|L(\widehat{f}_{\mathcal{V}_{m},\mathcal{D}_{\text{train}}},\bm{Z})|^{3}\mid\mathcal{D}_{\text{train}}]<\infty.$
Then Algorithm 2 with DA-MCS-plug2 yields
$\liminf_{n\to\infty}\mathbb{P}(\Theta^{*}\subseteq\widehat{\Theta}\mid\mathcal{D}_{\text{train}},\mathcal{V}_{1},\dots,\mathcal{V}_{M})\geq 1-\alpha.$
The condition on the third moment of the loss is mild and typically satisfied in practice.
@@ 3.5 MCS Ensembles
Usually, an MCS contains more than one model. In such cases, there is a set of good models whose performance is effectively indistinguishable given the limited amount of observations. In such cases, it is natural to take an ensemble approach. This may have the additional benefit of variance reduction and improved predictive performance, as is well-known from the literature on model averaging (Hoeting et al.,, 1999; Burnham and Anderson,, 2002). In particular, we propose to define the MCS mixture model
@@
For regression problems, we adapt the estimating equation approach of Nagler and Vatter, (2024) to the MCS mixture model. Specifically, let $\psi_{\beta}$ be a function such that $\mathds{E}[\psi_{\beta}(Y)\mid\bm{X}=\bm{x}]=0$ if and only if $\beta$ is the target of inference. Examples are $\psi_{\beta}(y)=y-\beta$ for the conditional mean $\beta(\bm{x})=\mathds{E}[Y\mid\bm{X}=\bm{x}]$ or $\psi_{\beta}(y)=\mathds{1}\{y<\beta(\bm{x})\}-\tau$ for the conditional $\tau$-quantile. We now approximate the conditional expectation by the fitted model, i.e., we solve
@@
with $w_{\widehat{\Theta}}(\bm{x},y)=\widehat{f}_{Y}(y)\sum_{\mathcal{V}\in\widehat{\Theta}}\widehat{c}_{\mathcal{V},\mathcal{D}}(\widehat{F}_{Y}(\bm{x}),\widehat{F}_{Y}(y))$, a weight that is the product of the copula density and the marginal density of $Y$. This approach is convenient because it puts no restrictions on the structure of the vine copula, facilitates model averaging, and can be used for all sorts of conditional quantities. We refer to Section 5 and Appendix C for additional discussions and computational details.
@@ 4 EXPERIMENTS
Table 1: Average negative log-likelihood (standard error) on test data for density estimation tasks. Best results in bold.
@@ 5 Setup
We use five data sets from the UCI repository (Dua and Graff,, 2017), Energy Efficiency ($n=768$, $p=8$), Concrete Compressive Strength ($n=1030$, $p=8$), Airfoil Self-Noise ($n=1503$, $p=5$), Wine quality ($n=4898$, $p=12$), Combined Cycle Power Plant (Ccpp) ($n=9568$, $p=5$), as well as the California Housing dataset ($n=20640$, $p=8$) (Pace and Barry,, 1997). All have a continuous label variable, a moderate number of features, and are commonly used for benchmarking density estimation and tabular regression models. For all datasets, we standardize the features to have mean zero and unit variance, randomly split the data into 80% training, and 20% test sets. For methods requiring a validation set, we further split the training set into 75% training and 25% validation.
@@
All experiments were run on a standard desktop machine. A Python package implementing the proposed methods as well as notebooks to fully reproduce the experiments is available in the supplementary materials and described in Appendix C. It relies mostly on the pyvinecopulib package (Vatter and Nagler,, 2025) for vine copula modeling. More specifically, for the marginals, we employ local-likelihood kernel density estimation with a bandwidth selection methodology similar to that of Sheather and Jones, (1991), with bias and variance expressions adapted for higher-order polynomials and nearest neighbor bandwidths. As for the pair-copulas, we use the nonparametric TLL estimator, which combines state-of-the-art performance with computational efficiency (Nagler et al.,, 2017). For MCS inference, we rely on a Python translation of the R (R Core Team,, 2025) script gracefully provided by the authors of Kim and Ramdas, (2025), improved to scale as $O(Mn)$ instead of $O(M^{2}n)$. All experiments were repeated 10 times with different random seeds. Averages and standard errors reported are over these replications, the latter in parentheses.
@@ 5.1 Density estimation
In this experiment, the target variable is stacked with the features, so that the problem becomes one of density estimation in $p+1$ dimensions. We compare the following methods:
  • Dissmann: A first benchmark method using the heuristic of Dissmann et al., (2013), based on absolute Kendall’s $\tau$.
  • Kraus: A second benchmark method using the heuristic of Kraus and Czado, 2017b , based on a combination of Kendall’s $\tau$ and $p$-values of tests for the simplifying assumption.
  • RS-B (M): Our random search Algorithm 1 with $M$ candidates, using the best model on the validation set with $L(f,\bm{Z})=-\log f(\bm{Z})$.
  • RS-E (M): An ensemble over the MCS obtained by Algorithm 2 with $M$ candidates with the same loss as RS-B and $\alpha=0.05$ (see Section 3.5).
In Table 1, we show the negative log-likelihood on test data. Except for the wine data set, both of our random search approaches outperform the benchmarks, irrespective of the number of candidates $M$. The performance improves with $M$, as expected, and our RS-B (500) method is the best in five out of six data sets, albeit by a small margin in some cases. The results also explain why the data set Energy is not included in Figure 1: the performance gain over the benchmark is simply too large to be displayed on the same scale. Given that model selection is based on the same loss as the evaluation, it is not surprising that RS-B outperforms RS-E, where emsembling may trade-off some accuracy for variance reduction.
@@
A key advantage of the MCS approach is that it enables assessing whether a benchmark is statistically indistinguishable from the best candidates using only the training data. Figure 3 reports the frequency with which Dissmann belongs to the $95\%$-MCS across datasets, which closely mirrors the results in Table 1. Notably, for the only dataset where Dissmann is the best-performing method (Wine), it is included in the MCS in every replication. In practice, this provides a simple check: the benchmark can be retained when it is competitive, without risking performance loss when it is clearly inferior.
@@ 5.2 Mean and median regression
  • Dissmann: The same heuristic as above, albeit combined with Nagler and Vatter, (2024) to turn the joint density estimate into a conditional mean or median regression function (see Section 3.5).
  • Kraus: The method of Kraus and Czado, 2017a , which uses the fact that D-Vines admit closed-form conditional distributions.
  • RS-B (M): Our random search Algorithm 1 with $M$ candidates, using the best model on the validation set with $L(f,\bm{Z})=-\log f(Y|\bm{X})$, turned into a regression function as for Dissmann.
  • RS-E (M): Same, albeit for Algorithm 2 with the MCS using $\alpha=0.05$, and Nagler and Vatter, (2024) adapted to ensembles as described in Section 3.5.
In Table 2, we show the root mean squared error (RMSE) on test data for mean regression. Similarly to the density estimation experiment, random search methods outperform the benchmarks in most cases, with performance improving with $M$. In this case however, RS-E (500) is the best method in four out of six datasets, with the RMSE of RS-E (M) being consistently lower than that of RS-B (M) for almost all $M$ and datasets. The exact same pattern is observed for median regression when evaluated via mean absolute error (MAE); see Appendix D for details. This suggests that ensembling via the MCS is particularly beneficial in regression-type problems, where the model selection is not directly optimized for the evaluation metric.
@@ 5.3 Probabilistic forecasting
In this experiment, we compare the same methods as in the regression experiment. However, we evaluate the quality of the full predictive distribution via the continuous ranked probability score (CRPS) (Gneiting and Katzfuss,, 2014), a popular proper scoring rule for probabilistic forecasting. Specifically, we use the fact that the methods from the previous experiment yield quantiles at all levels, and compute the CRPS via
@@
with $\rho_{\tau}(u)=u(\tau-\mathds{1}\{u<0\})$ the check function. Results are shown in Table 3. Outperformance of our random search methods over the benchmarks is even more pronounced in this experiment, and ensembles via the MCS are again particularly effective.
@@ 5.4 Runtime
In Figure 4, we show the relative CPU time of our random search methods compared to Dissmann on the Concrete data set. For both RS-B and RS-E, the computational cost is mostly determined by the number of candidates $M$. As expected, the CPU time grows roughly linearly with $M$ in training, irrespective of the random search method or the task. On the other hand, there is no increase in inference time for RS-B, since only one model is used for prediction. For RS-E, the inference time grows with $M$, since some fraction of the candidates are retained in the MCS and used for prediction.
While the computational cost of our random search methods is higher than that of Dissmann, it should be mentioned that the latter is almost negligible for real datasets of this scale. For instance, training and inference average respectively 0.357 and 0.006 seconds on a single CPU core for the Concrete data set. Additionally, the random search methods are easily parallelizable, since candidates can be fitted independently. See Appendix D for additional runtime results on other data sets.

### v2

@@ 3.3 Vine Confidence Sets
On any given data set, we may want to assess whether the model found by Algorithm˜1 is better than the benchmark heuristic of (Dissmann2013). When this is not the case, we might prefer the benchmark model for qualitative reasons such as ease of interpretability or easier comparison to previous analyses. And even if the benchmark is outperformed, multiple candidates from the random search may be statistically indistinguishable. We tackle this problem via model confidence sets (MCS) (hansen2011model); roughly speaking, subsets of the candidates that contain the “best ones” with high probability.
Define the set of optimal models according to expected out-of-sample loss as
$\Theta^{*}=\operatorname*{arg\,min}_{\mathcal{V}\in\Theta}\mathds{E}[L(\widehat{f}_{\mathcal{V},\mathcal{D}_{\text{train}}},\bm{Z})\mid\mathcal{D}_{\text{train}}],$
explicitly allowing for the possibility of multiple ones that are equally good. We emphasize that, in our setup, the goal is not to identify a fixed ground truth. We aim to find structures that lead to trained models with minimal out-of-sample loss. In this setting, the optimal set $\Theta^{*}$ is itself random and determined by the training data, estimation algorithm, and candidate set. In particular, we do not assume that the data was generated from a vine copula model. An MCS is then another random subset that contains each of the optimal models with high probability, as formalized in the following definition.
Definition 3.1 (MCS for Vines).
An $\alpha$-level MCS is a random set $\widehat{\Theta}\subseteq\Theta$ such that
@@
$\forall\,\mathcal{V}\in\Theta^{*}\colon\;\mathbb{P}(\mathcal{V}\in\widehat{\Theta}\mid\mathcal{D}_{\text{train}})\geq 1-\alpha.$
Such an MCS offers marginal guarantees, in the sense that the probabilistic statement applies to each optimal model separately, rather than to the set as a whole. An alternative concept is a uniform MCS requiring that $\Theta^{*}\subseteq\widehat{\Theta}$ with high probability. While this is a stronger guarantee, it is more difficult to achieve in practice, and may lead to unnecessarily large sets including many suboptimal models.
3.4 A Vine MCS Algorithm
Algorithm 2 MCS for Vines.
@@ 3.4 A Vine MCS Algorithm
3:Follow Steps 1-3 from Algorithm˜1.
4:Compute losses for $1\leq i\leq n_{\text{val}}$ and $1\leq m\leq M$
$L_{i,m}=L(\widehat{f}_{\mathcal{V}_{m},\mathcal{D}_{\text{train}}},\bm{Z}_{i}),\quad\bm{Z}_{i}\in\mathcal{D}_{\text{val}}.$
@@
Several methods for estimating MCSs have been proposed in the literature (hansen2011model; lei2020cross; mogstad2024inference; zhang2024winners; kissel2022black). They generally rely on discrete argmin inference, where the null hypothesis for the optimality of a given candidate $\mathcal{V}\in\Theta$ can be written as
$H_{0,\mathcal{V}}\colon\;\mathds{E}[L(\widehat{f}_{\mathcal{V},\mathcal{D}_{\text{train}}},\bm{Z})]\leq\min_{\mathcal{W}\in\Theta}\mathds{E}[L(\widehat{f}_{\mathcal{W},\mathcal{D}_{\text{train}}},\bm{Z})].$
If $\widehat{T}_{\mathcal{V}}$ is a statistic satisfying $\widehat{T}_{\mathcal{V}}\to N(0,1)$ under $H_{0,\mathcal{V}}$, a natural MCS estimator obtains by inverting the test:
$\widehat{\Theta}=\{\mathcal{V}:\widehat{T}_{\mathcal{V}}\leq\Phi^{-1}(1-\alpha)\,\},$
where $\Phi$ is the standard normal distribution function. In this work, we use the DA-test of kim2025locally because of its computational simplicity and good empirical performance. A full description of the test can be found in Appendix˜A for completeness, along with a runtime comparison with alternative methods.
Algorithm˜2 describes how this MCS method fits into our vine structure learning pipeline. The algorithm has runtime complexity $O(Mnd^{2}+Mn)$, where the first term is from the first three steps (see the discussion in Section˜3.2), the second from the MCS procedure, and we refer to Section˜A.2 for this part of the complexity analysis. It also comes with the theoretical guarantee of Proposition˜3.2, whose proof is given in Appendix˜B.
@@
Suppose
$\limsup_{n\to\infty}\max_{\mathcal{V}\in\Theta}\mathds{E}[|L(\widehat{f}_{\mathcal{V},\mathcal{D}_{\text{train}}},\bm{Z})|^{3}\mid\mathcal{D}_{\text{train}}]<\infty.$
Then Algorithm˜2 yields
$\liminf_{n\to\infty}\min_{\mathcal{V}\in\Theta^{*}}\mathbb{P}(\mathcal{V}\in\widehat{\Theta}\mid\mathcal{D}_{\text{train}})\geq 1-\alpha.$
As is standard, the guarantee is asymptotic. It is also conditional on the training data, which determines the optimal set $\Theta^{*}$. The condition on the third moment of the loss is mild and typically satisfied in practice.
@@ 3.5 MCS Ensembles
Usually, an MCS contains more than one model. In such cases, there is a set of good models whose performance is effectively indistinguishable given the limited amount of observations. In such cases, it is natural to take an ensemble approach. This may have the additional benefit of variance reduction and improved predictive performance, as is well-known from the literature on model averaging (hoeting1999bayesian; burnham2002model). In particular, we propose to define the MCS mixture model
@@
For regression problems, we adapt the estimating equation approach of Nagler2018 to the MCS mixture model. Specifically, let $\psi_{\beta}$ be a function such that $\mathds{E}[\psi_{\beta}(Y)\mid\bm{X}=\bm{x}]=0$ if and only if $\beta$ is the target of inference. Examples are $\psi_{\beta}(y)=y-\beta$ for the conditional mean $\beta(\bm{x})=\mathds{E}[Y\mid\bm{X}=\bm{x}]$ or $\psi_{\beta}(y)=\mathds{1}\{y<\beta(\bm{x})\}-\tau$ for the conditional $\tau$-quantile. We now approximate the conditional expectation by the fitted model, i.e., we solve
@@
with
$w_{\widehat{\Theta}}(\bm{x},y)=\widehat{f}_{Y}(y)\sum_{\mathcal{V}\in\widehat{\Theta}}\widehat{c}_{\mathcal{V},\mathcal{D}}(\widehat{F}_{Y}(\bm{x}),\widehat{F}_{Y}(y)),$
a weight that is the product of the copula density and the marginal density of $Y$. This approach is convenient because it puts no restrictions on the structure of the vine copula, facilitates model averaging, and can be used for all sorts of conditional quantities. We refer to Section˜4.1 and Appendix˜C for additional discussions and computational details.
@@ 4 Experiments
Table 1: Average NLL (standard error) on test data for density estimation tasks. Best results in bold.
@@ 4.1 Setup
We use five data sets from the UCI repository (Dua:2019), Energy Efficiency ($n=768$, $p=8$), Concrete Compressive Strength ($n=1030$, $p=8$), Airfoil Self-Noise ($n=1503$, $p=5$), Wine quality ($n=4898$, $p=12$), Combined Cycle Power Plant (Ccpp) ($n=9568$, $p=5$), as well as the California Housing dataset ($n=20640$, $p=8$) (Pace1997). All have a continuous label variables and are commonly used for benchmarking density estimation and tabular regression models. The number of features is in the typical range for applications of vine copula models, see also the discussions in Section˜2.4 and Section˜5 for additional challenges in higher-dimensional settings. For all datasets, we standardize the features to have mean zero and unit variance, randomly split the data into 80% training, and 20% test sets. For methods requiring a validation set, we further split the training set into 75% training and 25% validation.
@@
All experiments were run on a standard desktop machine. A Python package implementing the proposed methods and scripts to reproduce the experiments are available in the supplementary materials and described in Appendix˜C. It relies mostly on the pyvinecopulib package (pyvinecopulib) for vine copula modeling. Marginal densities and distribution are estimated by kernel local-likelihood. As for the pair-copulas, we use the nonparametric TLL estimator, which combines state-of-the-art performance with computational efficiency (nagler2017nonparametric).
For MCS inference, we rely on a Python translation of the R (R2025) script gracefully provided by the authors of kim2025locally, improved to scale as $O(Mn)$ instead of $O(M^{2}n)$; see Section˜A.2 for details. All experiments were repeated 10 times with different random seeds, and we report the average performance and standard error across these replications.
@@ 4.2 Density Estimation
In this experiment, we assess the quality of density estimation by stacking the features and the label, so that the problem becomes $(p+1)$-dimensional. We compare the following methods:
  • Dissmann: A first benchmark method using the heuristic of Dissmann2013, based on absolute Kendall’s $\tau$.
  • Kraus: A second benchmark method using the heuristic of Kraus2017, based on a combination of Kendall’s $\tau$ and $p$-values of tests for the simplifying assumption.
  • RS-B (M): Our random search Algorithm˜1 with $M$ candidates, using the best model on the validation set with $L(f,\bm{Z})=-\log f(\bm{Z})$.
  • RS-E (M): An ensemble over the MCS obtained by Algorithm˜2 with $M$ candidates with the same loss as RS-B and $\alpha=0.05$ (see Section˜3.5).
Other approaches mentioned in Section˜2.4 are not included in the comparison because of their prohibitive computational cost or lack of general applicability.
In Table˜1, we show the average negative log-likelihood (NLL) on test data. Except for the wine data set, both of our random search approaches outperform the benchmarks, irrespective of the number of candidates $M$. The performance improves with $M$, as expected, and our RS-E (500) method is the best on all data sets, albeit the margin is small in in some cases. The results also explain why the data set Energy is not included in Figure˜1: the performance gain over the benchmark is simply too large to be displayed on the same scale.
@@
A key benefit of the MCS approach is that it enables assessing whether a benchmark is statistically indistinguishable from the best candidates using only the training data. Figure˜3 reports the frequency with which Dissmann belongs to the $95\%$-MCS across datasets, which closely mirrors the results in Table˜1. Notably, for the only dataset where Dissmann is better than RS-B (Wine), it is included in the MCS in every replication. In practice, this provides a simple check: the benchmark can be retained when it is competitive, without risking performance loss when it is clearly inferior.
@@ 4.3 Mean and Median Regression
  • Dissmann: The same heuristic as above, albeit combined with Nagler2018 to turn the joint density estimate into a conditional mean or median regression function (see Section˜3.5).
  • Kraus: The method of kraus2017d, which uses the fact that D-Vines admit closed-form conditional distributions.
  • RS-B (M): Our random search Algorithm˜1 with $M$ candidates, using the best model on the validation set with $L(f,\bm{Z})=-\log f(Y|\bm{X})$, turned into a regression function as for Dissmann.
  • RS-E (M): Same, albeit for Algorithm˜2 with the MCS using $\alpha=0.05$, and Nagler2018 adapted to ensembles as described in Section˜3.5.
In Table˜2, we show the root mean squared error (RMSE) on test data for mean regression. Similarly to the density estimation experiment, random search methods outperform the benchmarks in most cases, with performance improving with $M$. Again, RS-E (500) is the best method in five out of six datasets, with the RMSE of RS-E (M) being consistently lower than that of RS-B (M) for almost all $M$ and datasets. The exact same pattern is observed for median regression when evaluated via mean absolute error (MAE); see Appendix˜D for details. This suggests that ensembling via the MCS is particularly beneficial in regression-type problems, where the model selection is not directly optimized for the evaluation metric.
@@ 4.4 Probabilistic Forecasting
In this experiment, we compare the same methods as in the regression experiment. However, we evaluate the quality of the full predictive distribution via the continuous ranked probability score (CRPS) (gneiting2014probabilistic), a popular proper scoring rule for probabilistic forecasting. Specifically, we use the fact that the methods from the previous experiment yield quantiles at all levels, and compute the CRPS via
@@
with $\rho_{\tau}(u)=u(\tau-\mathds{1}\{u<0\})$ the check function. Results are shown in Table˜3. Outperformance of our random search methods over the benchmarks is even more pronounced in this experiment, and ensembles via the MCS are again particularly effective.
@@ 4.5 Runtime
In Figure˜4, we show the relative CPU time of our random search methods compared to Dissmann on the Concrete data set. As expected, the CPU time grows roughly linearly with $M$ in training, irrespective of the random search method or the task. On the other hand, there is no increase in inference time for RS-B, since only one model is used for prediction. For RS-E, the inference time grows with $M$, since some fraction of the candidates are retained in the MCS and used for prediction.
While the computational cost of the random search methods is higher than that of Dissmann, the total cost is still tiny for real datasets of this scale. For instance, training and inference average respectively 0.357 and 0.006 seconds on a single CPU core for the Concrete data set. Additionally, the random search methods are easily parallelizable, since candidates can be fitted independently. See Appendix˜D for additional runtime results on other data sets.